Opened 10 years ago

Closed 9 years ago

#212 closed flight processing (fixed)

VOC_09, flight day 311/2008, Chile

Reported by: benj Owned by: mggr
Priority: alpha 4 high Milestone: 2008 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by benj)

Data location: /data/larsenb1/scratch/arsf/vocals/vocals-H/voc09-31108 (symlink in ~arsf/arsf_data/2008/flight_data/vocals/VOC_09-2008_311_Chile)

Data arrived from ARSF via SATA disk H on 12th December 2008.

Scientific details: VOCALS cloud campaign

Priority:

PI: G. Vaughan

Sensors:

  • Eagle
  • Hawk

Quicklook

TBD.

Change History (14)

comment:1 Changed 10 years ago by benj

  • Summary changed from VOC_08, flight day 310/2008, Chile to VOC_08, flight day 311/2008, Chile

comment:2 Changed 10 years ago by benj

  • Summary changed from VOC_08, flight day 311/2008, Chile to VOC_09, flight day 311/2008, Chile

comment:3 Changed 10 years ago by benj

  • Description modified (diff)

comment:4 Changed 10 years ago by jatt

Have renamed Eagle files so all start with VNIR prefix.

comment:5 Changed 10 years ago by jatt

  • Owner changed from jatt to benj

Delivery directory created at ~arsf/arsf_data/2008/flight_data/vocals/VOC_09-2008_311_Chile/delivery/20090429/VOC09

comment:6 Changed 10 years ago by benj

  • Owner changed from benj to peland

Passed to Peter for delivery checking

comment:7 Changed 10 years ago by peland

Checked logfile, seems fine.

Checking L1:
Some line dropouts across all samples and bands in Eagle data, eg e311031b.bil line 557 (saturation?). Images and spectra seem fine elsewhere.
Similar but wider line dropouts in Hawk, eg h311011b.bil lines 203-6, also sample dropouts eg sample 167 is low in band 68, 215 and 275 are zero in band 118. These seem to be present in all images at the same sample and band numbers - bad CCD elements? h311021b.bil has a greenish cast (green=band 68). Otherwise, images and spectra seem fine.
Apologies if this is all old hat - it's my first delivery check! 8)

comment:8 Changed 10 years ago by benj

  • Owner changed from peland to jatt

Line dropouts are expected - largely down to the sensor just losing lines periodically, as far as I'm aware. The sample dropouts are, as you guessed, bad CCD pixels - we have a ticket for re-doing the bad pixel file (#111), we just haven't done it yet. The greenish colour is a little odder, but the spectra look OK to me - I suspect it's just down to how ENVI's scaled it for some reason.

Should be OK then anyway, passing back to Jane for dispatch.

comment:9 Changed 10 years ago by jatt

All data delivered 18th May 2009.

comment:10 Changed 10 years ago by jatt

All flight lines are CM 69.

The pitch, roll and heading values used for this flight are:-

Sensor: Eagle

Flight Line Pitch Roll Heading
01-05 -0.2 -0.45 -0.5

Sensor: Hawk

Flight Line Pitch Roll Heading
01 -0.276 0.342 0.8
02 -0.28 0.34 0.79
03 -0.301 0.334 0.760
04 -0.335 0.334 0.803
05 -0.300 0.310 0.774

comment:11 Changed 10 years ago by mggr

Requested archiving by Sandra.

comment:12 Changed 10 years ago by mggr

  • Component changed from Processing: general to Archiving
  • Owner changed from jatt to mggr

Archived locally by DTG on 29/Jan/2009. Ready to go to NEODC.

comment:13 Changed 10 years ago by benj

Actually archived by DTG 15th Jun 2009. Still OK to go to NEODC.

comment:14 Changed 9 years ago by benj

  • Resolution set to fixed
  • Status changed from new to closed

Given to Jurgen (NEODC) on SATA disk "Linex backup 5" (sic) at PML on 15/01/2010

Note: See TracTickets for help on using tickets.